OpenStudy (anonymous):

Evaluate the limx>0 2-square root8/2-x and all divided by 1-square root 1-x

OpenStudy (sirm3d):

\[\Large \lim_{x\rightarrow 0} \frac{2-\sqrt{\frac{8}{2-x}}}{1-\sqrt{1-x}}\]

OpenStudy (anonymous):

yep

OpenStudy (sirm3d):

rationalize both sides of the fraction, one-time or one-at-a-time

OpenStudy (anonymous):

will i be rationalizing twice or once

OpenStudy (sirm3d):

if you can handle the one-time rationalizing of the fraction, go ahead. i'll be around to oversee your work

OpenStudy (anonymous):

ohk... let me work on it

OpenStudy (anonymous):